### v1

@@ 3.1 Causes
Misinformation. Misinformation has long been a significant concern in the modern digital age Shu et al. (2017); Zubiaga et al. (2018); Kumar and Shah (2018); Meel and Vishwakarma (2020); Fung et al. (2022); Wang et al. (2023b). The emergence of RAG introduces a novel approach to incorporating external documents to enhance the generation quality of LLMs. While this method has the potential to enrich content with diverse knowledge sources, it also poses the risk of including documents containing misinformation, such as fake news Chen et al. (2023b). Moreover, there have been instances where AI technologies have been employed to create or propagate misinformation Zhou et al. (2023c); Vergho et al. (2024); Weidinger et al. (2021). The advanced generative capabilities of LLMs exacerbate this issue, leading to an increase in misinformation generated by these systems. This trend is concerning, as it not only contributes to the spread of false information but also challenges detecting misinformation generated by LLMs Chen and Shu (2023b); Menczer et al. (2023); Barrett et al. (2023); Bengio et al. (2023); Wang et al. (2023c); Solaiman et al. (2023); Weidinger et al. (2023); Ferrara (2023); Goldstein et al. (2023).
Outdated Information. In addition to the challenge of misinformation, it is important to recognize that facts can evolve over time. The retrieved documents may contain updated and outdated information from the network simultaneously, leading to conflicts between these documents Chen et al. (2021); Liska et al. (2022); Zhang and Choi (2021); Kasai et al. (2022).
@@ 3.2 Analysis of Model Behaviors
Performance Impact. Previous research empirically demonstrates that the performance of a pre-trained language model can be significantly influenced by the presence of misinformation Zhang and Choi (2021) or outdated information Du et al. (2022b) within a specific context. In recent studies, Pan et al. (2023a) introduce a misinformation attack strategy involving the creation of a fabricated version of Wikipedia articles, which is subsequently inserted into the authentic Wikipedia corpus. Their research findings reveal that existing language models are susceptible to misinformation attacks, irrespective of whether the fake articles are manually crafted or generated by models. To gain a deeper understanding of how LLMs behave when encountering contradictory contexts, Chen et al. (2022) primarily conduct experiments using Fusion-in-Decoder on the NQ-Open Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017). They find that inconsistencies across knowledge sources exert a minimal effect on the confidence levels of models. These models exhibit a tendency to favor context directly pertinent to the query and context that aligns with the model’s inherent parametric knowledge. Xie et al. (2023) conduct experiments on both closed-source LLMs and open-source LLMs in PopQA Mallen et al. (2022) and StrategyQA Geva et al. (2021). The results obtained are in line with those of Chen et al. (2022), indicating that LLMs exhibit a significant bias to evidence that aligns with the model’s parametric memory. They also find that LLMs exhibit a predisposition towards emphasizing information related to entities of higher popularity and answers that are corroborated by a larger volume of documents within the given context. Moreover, these models demonstrate a significant sensitivity to the order in which data is introduced. Jin et al. (2024a) discover that as the number of conflicting hops increases, LLMs encounter increased challenges in reasoning.
@@
Detection Ability. In addition to assessing the performance of LLMs when confronted with contradictory contexts, several studies also investigate their capacity to identify such contradictions. Zheng et al. (2022) examine the performance of various models including BERT, RoBERTa, and ERNIE in detecting the contradiction within Chinese conversations. Their experimental findings reveal that identifying contradictory statements within a conversation is a significant challenge for these models. Li et al. (2023a) analyse the performance of GPT-4, ChatGPT, PaLM-2, and Llama 2 in identifying contradictory documents within news articles Hermann et al. (2015), stories Kočiský et al. (2018), and wikipedia Merity et al. (2017). The authors find that the average detection accuracy is subpar. The study also finds that LLMs face specific challenges when addressing certain types of contradictions, particularly those involving subjective emotions or perspectives. Additionally, the length of documents and the variety of self-contradictions have a minor influence on the detection performance. Wan et al. (2024) investigate the text features that affect LLMs’ assessment of document credibility when faced with conflicting information. They discover that existing models heavily prioritize the relevance of a document to the query but often overlook stylistic features that humans consider important, such as the presence of scientific references or a neutral tone in the text. Jin et al. (2024a) discover that LLMs encounter difficulty in distinguishing truthful information from misinformation. In addition, they find that LLMs favor evidence that appears most frequently within the context, and demonstrate confirmation bias for external information aligning with their internal memory.
@@
Remarks. When encountering conflict within a given context, the exhibited knowledge of the LLMs is significantly influenced. However, determining how the model responds to various contextual nuances remains an area requiring further exploration. While different models may share certain commonalities, disparities in behavior arise due to variations in their training data. Moreover, as the model’s knowledge is totally derived from textual information, its approach to discerning misinformation differs significantly from that of humans.
@@ 3.3 Solutions
Query Augmentation. Weller et al. (2022) explore a query augmentation technique that prompts GPT-3 to formulate new questions derived from the original inquiry. They then assess the confidence for each question’s answer by referencing the corresponding passages retrieved. Based on the confidence, they decide whether to rely on the original question’s prediction or aggregate predictions from the augmented questions with high confidence scores.
Remarks. Currently, strategies for addressing inter-context conflicts primarily involve relying on model knowledge or leveraging external knowledge such as retrieved documents. Moreover, augmenting LLM capabilities with external tools has emerged as a novel paradigm. Exploring the use of external tools to support LLMs in resolving inter-context conflicts is a promising approach. In addition, devising a unified and efficient approach to handle various conflict types remains a formidable challenge.
@@ 4.1 Causes
Remarks. Intra-memory conflicts in LLMs arise from three distinct causes that occur at different stages. Among these causes, training corpus bias stands out as the fundamental catalyst for these conflicts. Incongruities of knowledge in the training dataset result in inconsistencies within the knowledge encoded within the model’s parameters. Additionally, the decoding strategy indirectly contributes to exacerbating these conflicts. The inherent randomness of the sampling process during inference amplifies the inconsistencies in the model’s responses. Knowledge editing, which aims to post-update the model’s knowledge, can inadvertently introduce conflicting information into the LLM’s memory.
@@ 4.2 Analysis of Model Behaviors
Cross-lingual Inconsistency. The universality of true knowledge transcends surface form variations Ohmer et al. (2023), a principle that should ideally apply to LLMs. However, LLMs maintain distinct knowledge sets for different languages, leading to inconsistencies Ji et al. (2023). Wang et al. (2023e) investigate the challenges LLMs face in extending edited knowledge across languages, suggesting that knowledge related to different languages is stored separately within the model parameters. Qi et al. (2023) propose a metric named RankC for evaluating the cross-lingual consistency of LLMs’ factual knowledge. They employ this metric for analyzing multiple models and reveal a pronounced language dependence in the knowledge stored by LLMs, with no observed improvement in cross-lingual consistency with increased model size.
@@ 4.3.1 Improving Consistency
Output Ensemble. Mitchell et al. (2022) propose a method to mitigate the inconsistency of language models by leveraging a two-model architecture, involving the utilization of a base model responsible for generating a set of potential answers, followed by a relation model that evaluates the logical coherence among these answers. The final answer is selected by considering both the base model’s and the relation model’s beliefs. Zhao et al. (2023b) introduce a method to detect whether a question may cause inconsistency for LLMs. Specifically, they first use LLMs to rephrase the original question and obtain corresponding answers. They then cluster these answers and examine the divergence. The detection is determined based on the divergence level.
@@ 5 Challenges and Future Directions
Knowledge Conflicts in the Wild. Currently, the creation of knowledge conflicts predominantly relies on the artificial generation of incorrect or misleading information. In the real world, one of the most common situations where knowledge conflicts arise is in RALMs (Retrieval-Augmented Language Models), where conflicts are present in documents retrieved by the retrieval module directly from the Web. Current analysis approaches exist a gap in the experimental setup of knowledge conflict, suggesting that findings from those environments Xie et al. (2023); Wang et al. (2023g) might not easily transfer to practical applications. Recent studies have begun to investigate the scenario in the wild by curating conflicting documents based on actual search results from Google for open-ended questions Wan et al. (2024). Looking ahead, there is a growing interest in more research that assesses how well LLMs perform in real-world scenarios rather than artificially created conflicts, to better understand their capabilities.
Solution at a Finer Resolution. Currently, there is no one-size-fits-all solution to knowledge conflict due to its inherent complexity. Existing either assume a prior Shi et al. (2023b) or focus on a subclass of conflict Wang et al. (2023g). We believe that addressing this issue requires a more fine-grained approach, taking into account several factors. Firstly, the nature of the user’s query plays a crucial role. Subjective or debatable questions naturally lead to conflicts as they may have multiple valid answers Bjerva et al. (2020); Wan et al. (2024). Secondly, the source of conflicting information can vary, including misinformation, outdated facts, or partially correct data Uscinski and Butler (2013); Guo et al. (2022). Lastly, it is also important to consider user expectations, such as whether they prefer a single definitive answer from the LLM or are open to multiple perspectives Floridi (2023). Given these considerations, future solutions to mitigate knowledge conflicts must delve into these nuances, recognizing that knowledge conflict encompasses a spectrum of problems with diverse causes, manifestations, and potential resolutions. Collaboration between NLP and HCI researchers is appreciated for conducting thorough investigations and developing effective solutions.
@@
Interplay among the Conflicts. Current research in the knowledge conflict of LLMs primarily concentrates on investigating conflicts of a singular type  Wang et al. (2023g); Chen et al. (2022); Li et al. (2023d) or a joint study of inter-context and context-memory conflict Jin et al. (2024a); Xie et al. (2023). However, there is a notable dearth of research on the interaction between intra-memory conflict and the other two types of conflicts. Several papers have proposed the existence of knowledge circuits in LLMs Chughtai et al. (2024); Huang et al. (2023), which are closely related to intra-memory conflict. Addressing this gap is crucial for understanding the relationship between the internal knowledge inconsistency of the model and its behavior in response to the context. Moreover, exploring the synergistic effects of various conflict types could unveil underlying mechanisms of knowledge representation and processing in LLMs and help us to develop more robust and accurate LLMs in practice.
@@
Explainability. Most recent work analyzed LLMs’ behaviors amidst knowledge conflicts at the output level Xie et al. (2023); Wang et al. (2023g). While some studies have observed and explored the model’s confidence in its output, i.e., logits Xu et al. (2023); Jin et al. (2024a), there has been less focus on the internal mechanism of the model, like specific attention heads or neuron activations during conflicts. This gap highlights a need for more microscopic examinations to better comprehend how models decide when encounter conflicts. A recent study conducted by Jin et al. (2024b) advances this by investigating the interpretability of LLMs through information flow analysis, pinpointing pivotal points for conflict mitigation. They discover there are some attention heads with opposite effects in the later layers, where memory heads can recall knowledge from internal memory, and context heads can retrieve knowledge from the external context. Inspired by this, a novel technique named Pruning Head via Path Patching (PH3) is introduced to resolve conflicts efficiently without updating model parameters.
@@
Multilinguality. To date, research on knowledge conflict has primarily focused on the English language. Future studies could expand in two directions. Firstly, by examining LLMs to address knowledge conflicts in non-English prompts, leveraging the many advanced non-English models available (e.g., GLM Zeng et al. (2022) for Chinese) or LLMs with multilingual capability (e.g., GPT-4 Achiam et al. (2024)) and noting differences from English to account for unique language characteristics. Secondly, addressing inter-context conflict where multiple documents in different languages might be retrieved, potentially involving cross-language knowledge conflicts. Solutions could include employing translation systems Dementieva and Panchenko (2021) or, for low-resource languages, leveraging high-resource language evidence or employing knowledge distillation techniques.
@@
Multimodality. Current research on knowledge conflicts has mainly focused on single text modality, leaving the study of these conflicts in multimodal contexts as a promising area for future exploration. As LLMs evolve to process information across various formats—text, images Alayrac et al. (2022); Li et al. (2023b), video Ju et al. (2022); Zhang et al. (2023b), audio and speech Borsos et al. (2023); Wu et al. (2023)—the potential for conflicts escalates in growing complexity. For instance, textual documents might clash with visual data, or the tone of an audio clip might contradict the content of an accompanying caption. Future research on multimodal knowledge conflicts could focus on crafting advanced LLMs skilled in cross-modal reasoning and conflict resolution across diverse data types. This effort necessitates the enhancement of models’ capabilities to navigate the complex dynamics between different modalities and the development of targeted datasets for effective training and evaluation. Additionally, exploring how users perceive and manage multimodal conflicts, such as discrepancies between text and images, will offer valuable insights into improving LLMs for better human interaction.
@@ 6 Conclusion
Through this survey, we have extensively investigated knowledge conflicts, shedding light on their categorization, causes, how LLMs respond to these conflicts, and possible solutions. Our findings reveal that knowledge conflict is a multifaceted issue, with a model’s behavior being closely tied to the particular type of conflicting knowledge. Besides, there appears to be a more complex interplay among the three types of conflicts. Furthermore, we observe that existing solutions primarily address artificially constructed scenarios, neglecting the subtleties of conflicts by relying on assumed priors and thus sacrificing granularity and breadth. Given the growing use of retrieval-augmented language models (RALMs), we anticipate that knowledge conflicts faced by LLMs will only increase in complexity, underscoring the need for more comprehensive research in this area.
@@
No ethical considerations were necessary for this survey.

### v2

@@ 3.1 Causes
Misinformation. Misinformation has long been a significant concern in the modern digital age Shu et al. (2017); Zubiaga et al. (2018); Kumar and Shah (2018); Meel and Vishwakarma (2020); Fung et al. (2022); Wang et al. (2023b). The emergence of RAG incorporates external documents to enhance the generation quality of LLMs. While RAG has the potential to enrich content with diverse knowledge sources, it also poses the risk of including documents containing misinformation, such as fake news Chen et al. (2023b). Moreover, there have been instances where AI technologies are employed to create or propagate misinformation Weidinger et al. (2021); Zhou et al. (2023c); Vergho et al. (2024). The advanced generative capabilities of LLMs exacerbate this issue, leading to an increase in misinformation generated by these systems. This trend is concerning, as it not only contributes to the spread of false information but also challenges detecting misinformation generated by LLMs Chen and Shu (2023b); Menczer et al. (2023); Barrett et al. (2023); Bengio et al. (2023); Wang et al. (2023c); Solaiman et al. (2023); Weidinger et al. (2023); Ferrara (2023); Goldstein et al. (2023).
Outdated Information. In addition to the challenge of misinformation, it is important to recognize that facts can evolve. The retrieved documents may contain updated and outdated information from the network simultaneously, leading to conflicts between these documents Chen et al. (2021); Liska et al. (2022); Zhang and Choi (2021); Kasai et al. (2022); Schlichtkrull et al. (2023).
@@ 3.2 Analysis of Model Behaviors
Performance Impact. Previous research empirically demonstrates that the performance of a pre-trained language model can be significantly influenced by the presence of misinformation Zhang and Choi (2021) or outdated information Du et al. (2022b) within a specific context. In recent studies, Pan et al. (2023a) introduce a misinformation attack strategy involving the creation of a fabricated version of Wikipedia articles, which is subsequently inserted into the authentic Wikipedia corpus. Their research findings reveal that existing language models are susceptible to misinformation attacks, irrespective of whether the fake articles are manually crafted or generated by models. To gain a deeper understanding of how LLMs behave when encountering contradictory contexts, Chen et al. (2022) primarily conduct experiments using Fusion-in-Decoder on the NQ-Open Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017). They find that inconsistencies across knowledge sources exert a minimal effect on the confidence levels of models. These models tend to favor context directly pertinent to the query and context that aligns with the model’s inherent parametric knowledge. Xie et al. (2023) conduct experiments on both closed-source LLMs and open-source LLMs in PopQA Mallen et al. (2022) and StrategyQA Geva et al. (2021). The results obtained are in line with those of Chen et al. (2022), indicating that LLMs exhibit a significant bias to evidence that aligns with the model’s parametric memory. They also find that LLMs exhibit a predisposition towards emphasizing information related to entities of higher popularity and answers that are corroborated by a larger volume of documents within the given context. Moreover, these models demonstrate a significant sensitivity to the order in which data is introduced. Jin et al. (2024a) discover that as the number of conflicting hops increases, LLMs encounter increased challenges in reasoning.
@@
Detection Ability. In addition to assessing the performance of LLMs when confronted with contradictory contexts, several studies also investigate their capacity to identify such contradictions. Zheng et al. (2022) examine the performance of various models including BERT, RoBERTa, and ERNIE in detecting the contradiction within Chinese conversations. Experiments reveal that identifying contradictory statements within a conversation is a significant challenge for these models. Li et al. (2023a) analyse the performance of GPT-4, PaLM-2, and Llama 2 in identifying contradictory documents within news articles Hermann et al. (2015), stories Kočiský et al. (2018), and wikipedia Merity et al. (2017). The authors find that the average detection accuracy is subpar. The study also finds that LLMs face specific challenges when addressing certain types of contradictions, particularly those involving subjective emotions or perspectives. Additionally, the length of documents and the variety of self-contradictions have a minor influence on the detection performance. Wan et al. (2024) investigate the text features that affect LLMs’ assessment of document credibility when faced with conflicting information. They discover that existing models heavily prioritize the relevance of a document to the query but often overlook stylistic features that humans consider important, such as the presence of scientific references or a neutral tone in the text. Jin et al. (2024a) discover that LLMs encounter difficulty in distinguishing truthful information from misinformation. In addition, they find that LLMs favor evidence that appears most frequently within the context, and demonstrate confirmation bias for external information aligning with their internal memory.
@@
Remarks. When encountering conflict within a given context, the exhibited knowledge of the LLMs is significantly influenced. However, determining how the model responds to various contextual nuances remains an area requiring further exploration. While different models may share certain commonalities, disparities in behavior arise due to variations in their training data. Moreover, as the model’s knowledge is derived from textual information, its approach to discerning misinformation differs significantly from that of humans.
@@ 3.3 Solutions
Query Augmentation. Weller et al. (2022) explore a query augmentation technique that prompts GPT-3 to formulate new questions derived from the original inquiry. They then assess the confidence for each answer by referencing the corresponding passages retrieved. Based on the confidence, they decide whether to rely on the original question’s prediction or aggregate predictions from the augmented questions with high confidence scores.
Remarks. Strategies for addressing inter-context conflicts primarily rely on model knowledge or leveraging external knowledge such as retrieved documents. Recently, augmenting LLM with external tools has emerged as a new paradigm. Exploring the use of external tools to support LLMs in resolving inter-context conflicts could be a promising approach. On the other hand, devising unified and efficient methods to handle various conflict types remains a formidable challenge.
@@ 4.1 Causes
Remarks. Intra-memory conflicts in LLMs arise from three distinct causes that occur at different stages. Among these causes, training corpus bias stands out as the fundamental catalyst. Incongruities of knowledge in the training dataset result in inconsistencies within the knowledge encoded within the model’s parameters. Additionally, the decoding strategy indirectly contributes to exacerbating these conflicts. The inherent randomness of the sampling process during inference amplifies the inconsistencies in the model’s responses. Knowledge editing, which aims to post-update the model’s knowledge, can inadvertently introduce conflicting information into the LLM’s memory.
@@ 4.2 Analysis of Model Behaviors
Cross-lingual Inconsistency. The universality of true knowledge transcends surface form variations Ohmer et al. (2023), a principle that should ideally apply to LLMs. However, LLMs maintain distinct knowledge sets for different languages, leading to inconsistencies Ji et al. (2023); Xue et al. (2024). Wang et al. (2023f) investigate the challenges LLMs face in extending edited knowledge across languages, suggesting that knowledge related to different languages is stored separately within the model parameters. Qi et al. (2023) propose a metric named RankC for evaluating the cross-lingual consistency of LLMs’ factual knowledge. They employ this metric for analyzing multiple models and reveal a pronounced language dependence in the knowledge stored by LLMs, with no observed improvement in cross-lingual consistency with increased model size.
@@ 4.3.1 Improving Consistency
Output Ensemble. Mitchell et al. (2022) mitigate the inconsistency of language models by leveraging a two-model architecture, involving the utilization of a base model responsible for generating a set of potential answers, followed by a relation model that evaluates the logical coherence among these answers. The final answer is selected by considering both the base model’s and the relation model’s beliefs. Zhao et al. (2023b) introduce a method to detect whether a question may cause inconsistency for LLMs. Specifically, they first use LLMs to rephrase the original question and obtain corresponding answers. They then cluster these answers and examine the divergence. The detection is determined based on the divergence level.
@@ 5 Challenges and Future Directions
Knowledge Conflicts in the Wild. Currently, the creation of knowledge conflicts predominantly relies on the artificial generation of incorrect or misleading information. In the real world, one of the most common situations where knowledge conflicts arise is in RALMs (Retrieval-Augmented Language Models), where conflicts are present in documents retrieved by the retrieval module directly from the Web. Current analysis approaches exist a gap in the experimental setup of knowledge conflict, suggesting that findings from those environments Xie et al. (2023); Wang et al. (2023h) might not easily transfer to practical applications. Recent studies have begun to investigate the scenario in the wild by curating conflicting documents based on actual search results from Google for open-ended questions Wan et al. (2024). Looking ahead, there is a growing interest in more research that assesses how well LLMs perform in real-world scenarios rather than artificially created conflicts, to better understand their capabilities.
Solution at a Finer Resolution. Currently, there is no one-size-fits-all solution to knowledge conflict due to its inherent complexity. Existing either assume a prior Shi et al. (2023b) or focus on a subclass of conflict Wang et al. (2023h). We believe that addressing this issue requires a more fine-grained approach, taking into account several factors. Firstly, the nature of the user’s query plays a crucial role. Subjective or debatable questions naturally lead to conflicts as they may have multiple valid answers Bjerva et al. (2020); Wan et al. (2024). Secondly, the source of conflicting information can vary, including misinformation, outdated facts, or partially correct data Guo et al. (2022); Akhtar et al. (2023). Lastly, it is also important to consider user expectations, such as whether they prefer a single definitive answer from the LLM or are open to multiple perspectives Floridi (2023). Given these considerations, future solutions to mitigate knowledge conflicts must delve into these nuances, recognizing that knowledge conflict encompasses a spectrum of problems with diverse causes, manifestations, and potential resolutions. Collaboration between NLP and HCI researchers is appreciated for conducting thorough investigations and developing effective solutions.
@@
Interplay among the Conflicts. Current research in the knowledge conflict of LLMs primarily concentrates on investigating conflicts of a singular type  Wang et al. (2023h); Chen et al. (2022); Li et al. (2023d) or a joint study of inter-context and context-memory conflict Jin et al. (2024a); Xie et al. (2023). However, there is a notable dearth of research on the interaction between intra-memory conflict and the other two types of conflicts. Several papers have proposed the existence of knowledge circuits in LLMs Chughtai et al. (2024); Huang et al. (2023), which are closely related to the intra-memory conflict. Addressing this gap is crucial for understanding the relationship between the internal knowledge inconsistency of the model and its behavior in response to the context. Moreover, exploring the synergistic effects of various conflict types could unveil underlying mechanisms of knowledge representation and processing in LLMs and help us to develop more robust and accurate LLMs in practice.
@@
Explainability. Most recent work analyzed LLMs’ behaviors amidst knowledge conflicts at the output level Xie et al. (2023); Wang et al. (2023h). While some studies have observed and explored the model’s confidence in its output, i.e., logits Xu et al. (2023); Jin et al. (2024a); Wang et al. (2024b), there has been less focus on the internal mechanism of the model, like specific attention heads or neuron activations during conflicts. This gap highlights a need for more microscopic examinations to better comprehend how models decide when encounter conflicts. A recent study conducted by Jin et al. (2024b) advances this by investigating the interpretability of LLMs through information flow analysis, pinpointing pivotal points for conflict mitigation. They discover there are some attention heads with opposite effects in the later layers, where memory heads can recall knowledge from internal memory, and context heads can retrieve knowledge from the external context. Inspired by this, Pruning Head via Path Patching is introduced to resolve conflicts efficiently without updating model parameters.
@@
Multilinguality. To date, research on knowledge conflict has primarily focused on the English language. Future studies could expand in two directions. Firstly, by examining LLMs to address knowledge conflicts in non-English prompts, leveraging the many advanced non-English models available (e.g., GLM Zeng et al. (2022) for Chinese) or LLMs with multilingual capability (e.g., GPT-4 OpenAI (2024)) and noting differences from English to account for unique language characteristics. Secondly, addressing inter-context conflict where multiple documents in different languages might be retrieved, potentially involving cross-language knowledge conflicts. Solutions could include employing translation systems Dementieva and Panchenko (2021) or, for low-resource languages, leveraging high-resource language evidence (Xue et al., 2024) or employing knowledge distillation techniques.
@@
Multimodality. Current research on knowledge conflicts mainly focused on the text modality, leaving the study of these conflicts in multimodal contexts as a promising area for future exploration. As LLMs evolve to process information across various formats—images Alayrac et al. (2022); Li et al. (2023b), video Ju et al. (2022); Zhang et al. (2023b), and audio Borsos et al. (2023); Wu et al. (2023)—the potential for conflicts escalates in growing complexity. For instance, textual documents might clash with visual data, or the tone of an audio clip might contradict the content of an accompanying caption. Future research on multimodal knowledge conflicts could focus on crafting advanced LLMs skilled in cross-modal reasoning and conflict resolution across diverse data types. This effort necessitates the enhancement of models’ capabilities to navigate the complex dynamics between different modalities and the development of targeted datasets for effective training and evaluation. Additionally, exploring how users perceive and manage multimodal conflicts, such as discrepancies between text and images, will offer valuable insights into improving LLMs for better human interaction.
@@ 6 Conclusion
Through this survey, we have extensively investigated knowledge conflicts, shedding light on their categorization, causes, how LLMs respond to these conflicts, and possible solutions. Our findings reveal that knowledge conflict is a multifaceted issue, with a model’s behavior being closely tied to the particular type of conflicting knowledge. Besides, there appears to be a more complex interplay among the three types of conflicts. Furthermore, we observe that existing solutions primarily address artificially constructed scenarios, neglecting the subtleties of conflicts by relying on assumed priors and thus sacrificing granularity and breadth. Given the growing use of retrieval-augmented language models, we anticipate that knowledge conflicts faced by LLMs will only increase in complexity, underscoring the need for more comprehensive research.
@@
We mainly searched for papers published after 2021 using key terms including “knowledge conflict”, “knowledge inconsistency”, “knowledge gap”, inter alia, on Google Scholar and the ACL Anthology. After initially identifying these papers, the authors classified them through reading and continued to track related but overlooked papers using their citations. We also used Google Scholar to follow up on the latest papers citing these to avoid omissions.
For the quantitative analysis and comparison section (§ A.1), we did not conduct computational experiments but simply organized the result reported in other literature as is.
Appendix A Appendix
A.1 Quantitative Analysis and Comparison
In the context of a survey paper, while it is beneficial to include quantitative results and analyses concerning the impact of knowledge conflicts across various types of conflicts and the performance comparison of different mitigation strategies, it is not a strict requirement. We acknowledge the complexity and impracticality involved in conducting such quantitative experiments, particularly due to the use of disparate datasets in behavioral analyses, as well as the variance in the inherent knowledge of LLMs across different knowledge cut-off snapshots, as detailed in Table 1.
Moreover, establishing a “fair” comparison within the mitigation strategies segment poses its own set of challenges, given the diversity in objectives influenced by various assumed priors, such as the perceived accuracy of context or inherent knowledge, as discussed in the main text. Despite these intricacies, we opt to present quantitative results by compiling existing evaluations from a range of papers. It is imperative, however, to approach this analysis with caution, recognizing that original authors may have employed different datasets, LLMs variants, or even pursued contrasting objectives.
A.2 Quantitative Results on the Impact of Knowledge Conflicts
The comparison of quantitative results on the impact of the three types of knowledge conflicts is shown in Table 2. We pick the results of representative behavior analysis literature for comparison.
A.3 Quantitative Results on the Effectiveness of Mitigation Strategies
Reference Model Dataset Quantitative Results
Faithful to context
Shi et al. (2023a) Llama, OPT, GPT-Neo, and FLAN NQ-SWAP, MemoTrap, and NQ Their method improves GPT-Neo 20B by 54.4% on Memotrap and by 128% on NQ-SWAP where LLMs need to adhere to the given context.
Zhou et al. (2023d) ChatGPT and Llama2 MRC and Re-TACRED Compared to the zero-shot base prompts, their prompting method leads to a reduction of 32.2% for maintaining parametric knowledge for MRC and a 10.9% reduction for Re-TACRED on GPT-3.5. Similarly, on Llama2, there is a 39.4% reduction for MRC and a 57.3% reduction for Re-TACRED.
Discriminating misinformation
Hong et al. (2023) ChatGPT and FiD NQ and TQA The authors train a discriminator with about 80% F1 score and use it to improve models performance above 5%.
Pan et al. (2023b) ChatGPT NQ-1500 and CovidNews The author’s mitigation method improves the accuracy by more than 10%.
Disentangling sources
Wang et al. (2023h) ChatGPT KNOWLEDGE CONFLICT The authors’ method achieved over 80% F1 score on contextual knowledge conflict detection.
Table 3: Comparison of quantitative results on the effectiveness of various mitigation strategies w.r.t. their objectives.
The effectiveness of various mitigation strategies is quantitatively compared in Table 3. It is important to note that our analysis is limited to works addressing three predominant types of mitigating objectives within the context of memory conflicts. This selection is deliberate, as other types of mitigating objectives in different conflict categories do not yet have a substantial body of work that would allow for a meaningful cross-method comparison.